Please help with these problems. Show all steps
(Costs, 4 pts.) If the expenditure costs are $200 million, spent equally on 4 years (1, 2, 3, and 4) at a rate of 10%, what will be the NPV of this Rehabilitation project in year 0? 5. BONUS (Costs, 5 extra points). What will be the IRR for the Rehabilitation project?
Answer #1

$200 million spent equally in 4 years means per yer expenditure = $200 million / 4 = $50 million

So, here, A = $50 million

                i = 10%

                n = 4

NPV = A(P/A, i, n)

        = 50(P/A, 10%, 4)

        = 50(3.170)

        = $158.5 millio

Note: For IRR calculation Annual revenue data is require which is not mentioned in the question.